Two forces P & 2P act on the particle if the first is doubled & the second increased by 20N.The direction of the resultant remains unchanged.Find the value of P.

Grade:

5 Answers

Vikas TU
10 years ago

Net Force F1 = underroot(P2 + 4p2 + 4p2) = underroot(9p2)

F2 = underroot(4p2 + (2p + 20)2 + 4p.(2p+20)) = underroot(16p2 + 160p +400)

 

Now, F1 = F2

7p2 +160p +400 = 0

P = - 40/28 = -10/7 N

and

P = - 100 N

 

Where ''-'' minus sign indicates that the force is acting in opposite direction.

 

plz approve!

Bevkoof Singh
10 years ago

20

SOURAV MISHRA
10 years ago

since the direction of the resultant remains unchanged the angle which it makes with either of the forces remains unchanged.

so we can apply the result of similarity of triangles to get the result. but try to draw the vector triangles yourself.

P/2P = 2P/(2P + 20).

this gives P = 10N.

abc
6 years ago
u can refer SL Arora for this answer................anyways P=10.................................................................................................................................
ankit singh
3 years ago
Two forces equal to p and 2p Newton act on a particle if the first be doubled
Originally Answered: Two forces equal to P and 2P newton act on a particle. If the first be doubled and the second be increased by 20 newtons, the direction of the resultant is unaltered.